# Disentangling representations of retinal images with generative models

### v1

@@ Abstract
Retinal fundus images play a crucial role in the early detection of eye diseases and, using deep learning approaches, recent studies have even demonstrated their potential for detecting cardiovascular risk factors and neurological disorders. However, the impact of technical factors on these images can pose challenges for reliable AI applications in ophthalmology. For example, large fundus cohorts are often confounded by factors like camera type, image quality or illumination level, bearing the risk of learning shortcuts rather than the causal relationships behind the image generation process. Here, we introduce a novel population model for retinal fundus images that effectively disentangles patient attributes from camera effects, thus enabling controllable and highly realistic image generation. To achieve this, we propose a novel disentanglement loss based on distance correlation. Through qualitative and quantitative analyses, we demonstrate the effectiveness of this novel loss function in disentangling the learned subspaces. Our results show that our model provides a new perspective on the complex relationship between patient attributes and technical confounders in retinal fundus image generation.
Keywords: disentanglement, generative model, representation learning, spurious correlation, causality, latent subspaces, retinal fundus images
@@ 1 Introduction
Retinal fundus images are medical images that capture the back of the eye and show the retina, optic disc, macula, and blood vessels. Due to their nature as color photographs acquired through the pupil, they are inexpensive, non-invasive, and quick to obtain, making them available even in resource-constrained regions. Despite their high availability, fundus images can be used to detect the presence not only of diseases of the eye, but also cardiovascular risk factors or neurological disorders using deep learning (Poplin et al., 2018; Kim et al., 2020; Rim et al., 2020; Son et al., 2020; Sabanayagam et al., 2020; Xiao et al., 2021; Rim et al., 2021; Nusinovici et al., 2022; Cheung et al., 2022; Tseng et al., 2023).
@@
However, deep learning models require large datasets, and large medical imaging cohorts are often heterogeneous, with technical confounding being ubiquitous: for fundus images, variations in camera types, pupil dilation, image quality, and illumination levels can affect image generation, and biases due to hospital sites and patient study selection can lead to spurious feature correlations of these factors with biological variations in the data. Because deep learning models run the risk of learning shortcuts based on such spurious correlations (Geirhos et al., 2020) instead of specific patient-related image features (Fig. 1a), they may only perform well within the same distribution as the training data. For example, a retinal imaging dataset could be collected from two hospitals using cameras from different manufacturers (Fay et al., 2023). Camera A produces images appearing different in hue from camera B. The first hospital may treat predominantly Latin American patients, while the second hospital may treat predominantly Caucasians. In this scenario, it is simpler for a deep learning model to infer a patient’s ethnicity from a fundus image by recognizing hue rather than understanding the hidden causal relationship between ethnicity and phenotypic image features (Castro et al., 2020).
@@
One solution for dealing with confounders in data is subspace learning, which combines representation learning and disentanglement. The subspaces can be constructed from a causal graph using domain knowledge about what influences the image generation process (Castro et al., 2020). In our case, we assume a simplified causal model that splits the fundus image generation into patient characteristics and technical factors (Fig. 1a). We assume a shortcut connection due to spurious correlations between a patient attribute, such as ethnicity, and the camera as a technical factor (Fig. 1a, blue dashed arrow). To prevent shortcut learning, we aim to disentangle these factors of variation by encoding them into statistically independent latent subspaces that preferably match the ground truth generative factors (Bengio et al., 2013; Higgins et al., 2018).
@@
One common strategy for acquiring disentangled representations makes use of generative image models such as VAEs (Kingma and Welling, 2014). Generative models offer a valuable inductive bias for representation learning by reconstructing images from the latent space and can be extended to learn independent subspaces (Fig. 1b) in an unsupervised or supervised manner (Higgins et al., 2017; Klys et al., 2018; Tschannen et al., 2018). Previous methods primarily focused on achieving a disentangled latent space, often neglecting the simultaneous pursuit of high-quality image generation. However, when latent space disentanglement and image generation are addressed simultaneously, generative image models provide a built-in interpretation method through controllable feature changes in the image space (Fig. 1d).
@@
In this work, we close the gap of missing generative disentangled representation learning in conjunction with high-quality image generation in the domain of fundus images. Our primary objectives encompass three joint tasks: (1) disentangling learned representations, (2) preventing shortcut learning, and (3) enabling controllable high-resolution image generation.
@@
Figure 1: Disentangling representations of retinal images with generative models. In panel a, we present our conceptual causal graph outlining the image generation process (Fay et al., 2023). Within this framework, we break down retinal fundus image generation into patient-specific and technical factors of variation. We emphasize the potential shortcut for deep learning models when patient attributes exhibit spurious correlations with technical factors. In panel b, we introduce a strategy to circumvent shortcut learning by using a generative model as an independent subspace learner. Panel c shows our metric for disentanglement: an ideal confusion matrix between the available labels (columns) and the learned subspaces (rows). In panel d we demonstrate controllable fundus image generation by generating images from swapped ethnicity subspaces between the encoded original images in the first column.
@@ 2 Related work
Generative models such as VAEs are established methods for representation learning (Tschannen et al., 2018). However, it is difficult to use them for detailed high-resolution reconstructions, which are relevant in the field of retinal imaging, where fine details have a large impact on patient attribute or disease classification. Consequently, work on generative representation learning primarily focuses on learning low-dimensional latent spaces with specific properties rather than trying to generate accurate image reconstructions (Tschannen et al., 2018).
Generative adversarial networks (GANs), on the other hand, succeed in the generation of fine details in the image domain. Especially state-of-the art GANs like StyleGAN2 (Karras et al., 2020) are able to create high-fidelity and high-resolution images. When labels are available, conditional GANs (Mirza and Osindero, 2014) and their variants (Odena et al., 2017; Miyato and Koyama, 2018) are the most established extensions for class-conditional image generation. However, conditional GANs mainly focus on generating realistic images and do not focus on the independence of the latent space from the conditioned label.
InfoGAN (Chen et al., 2016) modifies the GAN objective to learn unsupervised disentangled representations by maximizing a lower bound of the mutual information (MI) between a subset of the latent variables and the observation. However, this extension does not guarantee that for complex image data, their induced factored latent codes are not dominated by shortcut features such as camera type. Furthermore, due to the unsupervised setting, disentangled features are analyzed post-hoc with human feedback, leading to potential non-identifiability (Locatello et al., 2019). Semi-StyleGAN (Nie et al., 2020), a special case of semi-supervised conditional GANs, extends the StyleGAN architecture to an InfoStyleGAN with an MI loss and addresses non-identifiability by adding a weakly supervised reconstruction term for partially available factors of variation.
@@
In the field of disentanglement, a common goal is to use generative models as independent subspace learners, necessitating the estimation and minimization of statistical dependence between subspaces. Mutual Information (MI) is a viable measure for quantifying the mutual dependence between two variables, effectively capturing nonlinear dependencies. However, estimating MI in high-dimensional spaces remains a challenge. Earlier MI estimation approaches include non-parametric binning (Darbellay and Vajda, 1999), kernel density estimation (Härdle et al., 2006), likelihood-ratio estimation (Suzuki et al., 2008) and K-nearest neighbor estimation (Kraskov et al., 2004). These estimators typically do not scale well with sample size or dimension (Belghazi et al., 2018). Also, the gradient of these methods is difficult to calculate, which makes them inapplicable to back-propagation frameworks for MI optimization.
Therefore, recent work has focused on estimating bounds to establish tractable and scalable MI objectives. Most of the proposed approaches are lower bounds (Poole et al., 2019; Belghazi et al., 2018), which are not applicable to MI minimization problems. There is limited work on estimating MI upper bounds, and most of them are restrictive in the sense that they require the conditional distribution to be known (Alemi et al., 2017; Poole et al., 2019; Cheng et al., 2020).
Distance measures between distributions, such as the MI, provide a theoretically sound measure, but often require density estimation. To address this, non-parametric kernel method such as maximum mean discrepancy (MMD) (Sejdinovic et al., 2013) offer a principled alternative. MMD simplifies distance computation by using mean kernel embeddings of features, and is therefore also used to learn invariant representations, as in the Variational Fair Autoencoder (VFAE) (Louizos et al., 2017). Another non-parametric alternative to MI estimation is the use of adversarial classifiers, which solve the invariant subspace problem by solving the trade-off between task performance and invariance through iterative minimax optimization (Ganin and Lempitsky, 2015; Ganin et al., 2016; Xie et al., 2017). However, adversarial training adds an unnecessary layer of complexity to the learning problem Moyer et al. (2018) that does not scale well with multiple subspaces.
@@
In medical imaging, disentangled representations are relevant to many research questions related to image synthesis, segmentation, registration, and causal- or federated learning, e.g., for disease decomposition, harmonization, or controllable synthesis. For a more comprehensive overview of the applications in medical imaging, we refer to Liu et al. (2022). For example, Fay et al. (2023) present a predictive approach to avoid shortcut learning for MRI data from different scanners. To minimize the dependencies between a task-specific factor (e.g. patient attribute) and a confounding correlated factor (e.g. scanner), they estimate a lower bound for MI. However, as mentioned earlier, a lower bound MI estimate is inconsistent to MI minimization tasks. In ICAM (Bass et al., 2020, 2021), the authors build on a VAE-GAN approach for image-to-image translation (Lee et al., 2018) to disentangle disease-relevant from disease-irrelevant features. However, disentanglement in ICAM requires longitudinal data, where patient images are available at different disease stages. In a similar approach, Ouyang et al. (2021) embed MRI images into independent anatomical and modality subspaces by image-to-image translation. Like ICAM, they have the data constraint that their method only works with images of the same patient from different modalities. In another attempt to explain classifier features, StylEx (Lang et al., 2021) integrates a classifier into StyleGAN training. They have a classifier-specific intermediate latent space and can visualize image features that are important for the classifier decision. Although the authors apply their approach to the retinal fundus domain, they do not address spurious correlation and shortcut learning.
@@ 3 Methods
This section introduces the proposed extensions to the generative model architecture and training constraints to disentangle the representations, prevent shortcut learning, and still enable controllable high-resolution image generation. First, we introduce the dataset, an encoder model and our disentanglement loss as preliminary steps towards the complete image generation pipeline.
@@ 3.1 Data
Figure 2: Image preprocessing and data distribution. In panel a, we outline our two-step preprocessing of fundus images. The first column displays original samples from the training set, while the second column exhibits images tightly cropped around the fundus circle. In the third column, we further cropped and flipped the images for consistency. The corresponding label distributions are visualized as histogram plots in panels b-d. Specifically, panel b illustrates the age group distribution across three classes, panel c showcases the ethnicity distribution, and panel d presents the camera distribution. For underrepresented categories, the number of images is shown next to the respective bars.
@@
The retinal fundus images in EyePACS are not standardized in aspect ratio and field of view (Fig. 2a, first column). To do so, we cropped them to a tightly centered circle (Müller et al., 2023) (Fig. 2a, middle column) and resized them to a resolution of $256\times 256$ pixels. Furthermore, to address evident and unwanted factors of variation, we standardized the images by masking them to a uniform visible area, as many images had missing regions at the top and bottom. Additionally, we horizontally flipped the images to ensure that all optic discs were positioned on the left side (Fig. 2a, last column).
Retinal fundus images from EyePACS come with extensive technical and patient-specific metadata. Here, we relied on age, ethnicity, and camera labels (Fig. 2b-d). The camera labels in EyePACS include some duplicate manufacturer names, which we merged into the 14 classes (Fig. 2d).
@@ 3.2 Encoder model for mapping factors of variation
We first considered the image encoding task by learning an encoder $f_{\theta}$ that maps images to a latent representation $f_{\theta}:\mathcal{X}\rightarrow\mathcal{W}$. We observed noisy image data $x\in\mathcal{X}$ obtained from an unknown image generation process $x=g(s)$, where $s=(s_{1},s_{2},\dots,s_{k})$ are the underlying factors of variation. Our goal was to find a mapping $f_{\theta}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{d}$ for a latent space $f_{\theta}(x)=w=(w_{1},w_{2},\dots,w_{k})$ such that each factor of variation $s_{i}$ can be recovered from the corresponding latent subspace $w_{i}\in\mathbb{R}^{d_{i}}$, $d=\Sigma_{i=0}^{k}d_{i}$, by a linear mapping $\hat{s}_{i}=C_{\psi_{i}}w_{i}$, but not from the other latent subspaces $w_{j}$ for $j\neq i$.
@@
In practice, we worked with retinal fundus images as our input domain $\mathcal{X}$, using a vector of available metadata as labels $y$. The goal was to encode the images into disentangled subspaces: $w=(w_{1},w_{2})$ for potentially highly correlated labels. To enforce attribute encoding in the subspaces, we defined a predictive model (Fig. 3) consisting of two parts: the feature encoder and the classification heads. The feature encoder $f_{\theta}$ maps input images $x\in\mathbb{R}^{m}$ to feature vectors $w\in\mathbb{R}^{d}$. Each feature vector was split into different subspaces $w_{1},\dots,w_{K}$ by a linear classification head $C_{\psi_{k}}$ attached to each subspace $k$. Thus, optimizing the encoder to map to attribute subspaces was equivalent to optimizing the cross-entropy loss for each subspace individually:

### v2

@@ Abstract
Retinal fundus images play a crucial role in the early detection of eye diseases. However, the impact of technical factors on these images can pose challenges for reliable AI applications in ophthalmology. For example, large fundus cohorts are often confounded by factors like camera type, bearing the risk of learning shortcuts rather than the causal relationships behind the image generation process. Here, we introduce a population model for retinal fundus images that effectively disentangles patient attributes from camera effects, enabling controllable and highly realistic image generation. To achieve this, we propose a disentanglement loss based on distance correlation. Through qualitative and quantitative analyses, we show that our models encode desired information in disentangled subspaces and enable controllable image generation based on the learned subspaces, demonstrating the effectiveness of our disentanglement loss. The project’s code is publicly available [^1] https://github.com/berenslab/disentangling-retinal-images.
Keywords: disentanglement, generative model, spurious correlation, causality, retinal fundus images
@@ 1 Introduction
Retinal fundus images are medical images that capture the back of the eye and show the retina, optic disc, macula, and blood vessels. Due to their nature as color photographs acquired through the pupil, they are inexpensive, non-invasive, and quick to obtain, making them available even in resource-constrained regions. Despite their high availability, fundus images can be used to detect the presence not only of eye diseases, but also cardiovascular risk factors or neurological disorders using deep learning (Poplin et al., 2018; Kim et al., 2020; Rim et al., 2020; Son et al., 2020; Sabanayagam et al., 2020; Xiao et al., 2021; Rim et al., 2021; Nusinovici et al., 2022; Cheung et al., 2022; Tseng et al., 2023).
@@
However, deep learning models require large datasets, and large medical imaging cohorts are often heterogeneous, with technical confounding being ubiquitous. For fundus images, variations in camera types, pupil dilation, image quality, and illumination levels can affect image generation. Additionally, biases due to different hospital sites and patient study selection can lead to spurious feature correlations of these factors with biological variations in the data. Deep learning models run the risk of learning shortcuts based on such spurious correlations (Geirhos et al., 2020), instead of specific patient-related image features (Fig. 1a). As a result, they may only perform well within the same distribution as the training data and latent space analyses may be misleading (Müller et al., 2022). For example, a retinal imaging dataset could be collected from two hospitals using cameras from different manufacturers (Fay et al., 2023). Camera A produces images appearing different in hue from camera B. The first hospital may treat predominantly Latin American patients, while the second hospital may treat predominantly Caucasians. In this scenario, it is simpler for a deep learning model to infer a patient’s ethnicity from a fundus image by recognizing hue rather than understanding the hidden causal relationship between ethnicity and phenotypic image features (Castro et al., 2020).
@@
One solution for dealing with confounders in data is subspace learning, which combines representation learning and disentanglement. Latent subspaces can be constructed from a causal graph using domain knowledge about what influences the image generation process (Castro et al., 2020). In our case, we assume a simplified causal model that splits the fundus image generation into patient characteristics and technical factors (Fig. 1a). We assume a shortcut connection due to spurious correlations between a patient attribute, such as ethnicity, and the camera as a technical factor (Fig. 1a, blue dashed arrow). To prevent shortcut learning, we aim to disentangle these factors of variation by encoding them into statistically independent latent subspaces (Bengio et al., 2013; Higgins et al., 2018).
@@
One common strategy for acquiring disentangled representations makes use of generative image models such as variational autoencoders (VAEs) (Kingma and Welling, 2014). Generative models offer a valuable inductive bias for representation learning by reconstructing images from the latent space and can be extended to learn independent subspaces (Fig. 1b) (Higgins et al., 2017; Klys et al., 2018; Tschannen et al., 2018). Previous methods primarily focused on achieving a disentangled latent space, often neglecting the simultaneous pursuit of high-quality image generation. However, when latent space disentanglement and image generation are addressed simultaneously, generative image models provide a built-in interpretation method through controllable feature changes in the image space (Fig. 1d).
@@
In this work, we close the gap of missing generative disentangled representation learning in conjunction with high-quality image generation in the domain of fundus images. Our contributions are threefold
  • 1.
    Disentangling latent subspaces in a state-of-the-art generative adversarial model with a disentanglement loss based on distance correlation.
  • 2.
    Developing a method that jointly optimizes for disentanglement and controllable high-resolution image generation end-to-end.
  • 3.
    Conducting extensive analysis on the effectiveness of our disentanglement loss using both quantitative and qualitative measures to evaluate disentanglement, image quality, and controllable image generation.
@@
Figure 1: Disentangling representations of retinal images with generative models. In panel a, we present our conceptual causal graph outlining the image generation process (Fay et al., 2023). Within this framework, we break down retinal fundus image generation into patient-specific and technical factors of variation and highlight a potential shortcut connection for deep learning models. In panel b, we introduce a strategy to prevent shortcut learning by using a generative model as an independent subspace learner. Panel c shows our metric for disentanglement: an ideal confusion matrix between the available labels (columns) and the learned subspaces (rows). Panel d shows controllable fundus image generation from swapped ethnicity subspaces.
@@ 2 Related work
The most commonly used generative models for representation learning are VAEs (Tschannen et al., 2018). However, VAEs produce blurry images that lack details, which are relevant in the field of retinal imaging, where fine details have a major impact on patient attribute or disease classification. Consequently, work on generative representation learning with VAEs focuses primarily on learning low-dimensional latent spaces with specific properties, rather than on generating accurate image reconstructions (Tschannen et al., 2018).
Generative adversarial networks (GANs), on the other hand, succeed in generating fine details in the image domain and can be adapted to representation learning by GAN-inversion (Ghosh et al., 2022). Especially state-of-the art GANs like StyleGAN2 (Karras et al., 2020) are able to create high-fidelity and high-resolution images. When labels are available, conditional GANs (Mirza and Osindero, 2014) and their variants (Odena et al., 2017; Miyato and Koyama, 2018) are the most established extensions for class-conditional image generation. However, conditional GANs mainly focus on generating realistic images and do not focus on the statistical independence between the latent space and the conditioned label.
InfoGAN (Chen et al., 2016; Paul et al., 2021) modifies the GAN objective to learn unsupervised disentangled representations by maximizing a lower bound of the mutual information (MI) between a subset of the latent variables and the observation. However, this extension does not guarantee that for complex image data their induced factored latent codes are not dominated by shortcut features such as camera type. Furthermore, due to the unsupervised setting, disentangled features are analyzed post-hoc with human feedback, leading to potential non-identifiability (Locatello et al., 2019), especially in highly specialized domains such as medical imaging. Semi-StyleGAN (Nie et al., 2020), closest to our work, is a special case of semi-supervised conditional GANs and extends the StyleGAN architecture to an InfoStyleGAN with a mutual information loss and addresses non-identifiability by adding a weakly supervised reconstruction term for partially available factors of variation. However, they do not address the problem of potential spurious correlations in the data and the resulting need for statistically independent latent factors.
Diffusion probabilistic models (DPMs) show remarkable success in image generation. However, since the latent variables of DPMs lack high-level semantic meaning, little attention has been paid to representation learning with DPMs. There are few works that combine disentangled representation learning with DPMs. In (Preechakul et al., 2022; Zhang et al., 2022), they explore DPMs for representation learning via autoencoding by jointly training an encoder for high-level semantics and a conditional DPM as a decoder for image reconstruction. In (Yang et al., 2023), they eliminate semantic linear classifiers and achieve and unsupervised disentanglement by decomposing the gradient fields of DPMs into sub-gradient fields for different generative factors. However, compared to the StyleGAN2 architecture used in this work, DPMs lack the ability to control scale-specific image generation, the learned representations may not be easy to interpret, and image generation speed is considerably slower.
@@
In the field of disentanglement, a common goal is to use generative models as independent subspace learners, necessitating the estimation and minimization of statistical dependence between subspaces. Mutual information (MI) is a viable measure for quantifying the mutual dependence between two variables, effectively capturing nonlinear dependencies. However, estimating MI in high-dimensional spaces remains a challenge. Therefore, recent work focused on estimating bounds to establish tractable and scalable MI objectives. Most of the proposed approaches are lower bounds (Poole et al., 2019; Belghazi et al., 2018), which are not applicable to MI minimization problems. There is limited work on estimating MI upper bounds, and most of them are restrictive in the sense that they require the conditional distribution to be known (Alemi et al., 2017; Poole et al., 2019; Cheng et al., 2020).
Distance measures between distributions, such as the MI, provide a theoretically sound measure, but often require density estimation. To address this, non-parametric kernel method such as maximum mean discrepancy (MMD) (Sejdinovic et al., 2013) offer a principled alternative. MMD simplifies distance computation by using mean kernel embeddings of features, and is therefore also used to learn invariant representations, as in the Variational Fair Autoencoder (VFAE) (Louizos et al., 2017). Another non-parametric alternative to MI estimation are adversarial classifiers, which solve the invariant subspace problem by solving the trade-off between task performance and invariance through iterative minimax optimization (Ganin and Lempitsky, 2015; Ganin et al., 2016; Xie et al., 2017). However, adversarial training adds an unnecessary layer of complexity to the learning problem (Moyer et al., 2018) that does not scale well with multiple subspaces. In this work, we use distance correlation (Székely et al., 2007), as another non-parametric measure that does not require any pre-defined kernels. Distance correlation is fast to evaluate and captures the linear and nonlinear dependence between two random vectors of arbitrary dimensions. It proves to be of versatile use in deep learning, ranging from measuring disentanglement to measuring similarity between networks (Liu et al., 2021; Zhen et al., 2022).
@@
In medical imaging, disentangled representations are relevant to many research questions related to image synthesis, segmentation, registration, and causal- or federated learning, e.g., for disease decomposition, harmonization, or controllable synthesis Liu et al. (2022). For example, Fay et al. (2023) present a predictive approach to avoid shortcut learning for MRI data from different scanners. To minimize the dependencies between a task-specific factor (e.g. patient attribute) and a confounding correlated factor (e.g. scanner), they estimate a lower bound for MI. However, as mentioned above, a lower bound MI estimate is inconsistent with MI minimization tasks. In ICAM (Bass et al., 2020, 2021), the authors build on a VAE-GAN approach for image-to-image translation (Lee et al., 2018) to disentangle disease-relevant from disease-irrelevant features. However, disentanglement in ICAM requires longitudinal data, where patient images are available at different disease stages. In a similar approach, Ouyang et al. (2021) embed MRI images into independent anatomical and modality subspaces by image-to-image translation. Like ICAM, they have the data constraint that their method only works with images of the same patient from different modalities. In another attempt to explain classifier features, StylEx (Lang et al., 2021) integrates a classifier into StyleGAN training. They have a classifier-specific intermediate latent space and can visualize image features that are important for the classifier decision. Although the authors apply their approach to the retinal fundus domain, they do not address spurious correlation and shortcut learning.
@@ 3 Methods
This section introduces the proposed extensions to the generative model and training constraints to disentangle the representations, prevent shortcut learning, and still enable controllable high-resolution image generation. First, we introduce the dataset, an encoder model and our disentanglement loss as preliminary steps towards the complete image generation pipeline.
@@ 3.1 Data
Figure 2: Image preprocessing and data distribution. In panel a, we outline our two-step preprocessing of fundus images. The first column displays original samples from the training set, while the second column shows images tightly cropped around the fundus circle. In the third column, we further cropped and flipped the images for consistency. The corresponding label distributions are visualized as histogram plots in panels b-d. Specifically, panel b illustrates the age group distribution across three classes, panel c shows the ethnicity distribution, and panel d presents the camera distribution.
@@
The retinal fundus images in EyePACS are not standardized in aspect ratio and field of view (Fig. 2a, first column). To do so, we cropped them to a tightly centered circle (Müller et al., 2023) and resized them to a resolution of $256\times 256$ pixels (Fig. 2a, middle column). Furthermore, to address evident and unwanted factors of variation, we standardized the images by masking them to a uniform visible area, as many images had missing regions at the top and bottom. Additionally, we horizontally flipped the images to ensure that all optic discs were positioned on the left side (Fig. 2a, last column).
Retinal fundus images from EyePACS come with extensive technical and patient-specific metadata. Here, we relied on age, ethnicity, and camera labels (Fig. 2b-d). The camera labels in EyePACS include some duplicate manufacturer names, which we merged into 14 classes (Fig. 2d).
@@ 3.2 Encoder model for mapping factors of variation
We first considered the image encoding task by learning an encoder $f_{\theta}$ that maps images to a latent representation $f_{\theta}:\mathcal{X}\rightarrow\mathcal{W}$. We observed noisy image data $x\in\mathcal{X}$ obtained from an unknown image generation process $x=g(s)$, where $s=(s_{1},s_{2},\dots,s_{K})$ are the underlying factors of variation. Our goal was to find a mapping $f_{\theta}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{d}$ for a latent space $f_{\theta}(x)=w=(w_{1},w_{2},\dots,w_{K})$ such that each factor of variation $s_{k}$ can be recovered from the corresponding latent subspace $w_{k}\in\mathbb{R}^{d_{k}}$, $d=\Sigma_{k=1}^{K}d_{k}$, by a linear mapping $\hat{s}_{k}=C_{\psi_{k}}w_{k}$, but not from the other latent subspaces $w_{l}$ for $l\neq k$.
@@
In practice, we worked with retinal fundus images as our input domain $\mathcal{X}$, using a vector of available metadata as labels $y$. The goal was to encode the images into disentangled subspaces: $w=(w_{1},w_{2})$ for potentially highly correlated labels. To enforce attribute encoding in the subspaces, we defined a predictive model (Fig. 3) consisting of two parts: the feature encoder and the classification heads. The feature encoder $f_{\theta}$ maps input images $x\in\mathbb{R}^{m}$ to a feature vector $w\in\mathbb{R}^{d}$. Each feature vector was split into two subspaces by a linear classification head $C_{\psi_{k}}$ attached to each subspace $k$. Thus, optimizing the encoder to map to attribute subspaces was equivalent to optimizing the cross-entropy loss for each subspace individually: